# UniVST: A Unified Framework for Training-free Localized Video Style Transfer

### v4

@@ 1 Introduction
Video stylization, a sub-area of video editing, is gaining popularity. It is about adding artistic styles to video content. Our review identifies two main video stylization methods. The first method uses the image or text to direct the style. StyleCrafter [11] personalizes video style with an extra training adapter; VideoBooth [8] uses the image for custom video creation prompts; while Diffutoon [12] focuses on cartoon coloring from a text prompt. The second method includes general video editing frameworks like AnyV2V [13] and BIVDiff [10] that can perform various tasks, including style transfer. These usually apply a consistent style across the whole video, which is good for thematic expression but sometimes lacking in precision for specific uses. For example, as shown in Fig. 1, in film, scenes might be shot normally and later stylized to match different environments. This enables varied scene changes from one video take, which is also useful in advertising for efficiently applying local style changes to existing videos. Therefore, selectively styling parts of a video while keeping others unchanged is an important academic and commercial task.
@@
We illustrate Fig. 1 and pinpoint three major limitations of current methods: 1. Lack of Fine-grained Control. While some existing approaches [6, 5] leverage localized text descriptions to preserve the overall style transfer and maintain the primary object’s style, the control is usually too coarse. This lack of precision can lead to the model failing to understand the nuances, potentially causing unintended style transfer effects on objects not explicitly mentioned in the text. 2. Content Fidelity vs. Style Richness. In the realm of video stylization, there is a delicate balance to be struck between content fidelity and style richness. An overemphasis on style richness can result in blurring and alter the original layout of the video, whereas an excessive focus on content fidelity might yield a stylized outcome that is indistinguishable from the original. 3. Temporal Inconsistency. Unlike image stylization, video stylization requires careful consideration of temporal coherence between frames. As in BIVDiff [10] and AnyV2V [13], simply extending image stylization techniques to video often results in frame inconsistencies, causing various flickering and artifacts.
@@
In this paper, we propose UniVST, a unified localized [^1] Here, “localized” indicates a particular part of the video. video style transfer framework that addresses the above issues in a training-free manner. The first step in achieving this is to obtain masks of the region specified by users, which allows for precise control during the subsequent stylization process. However, unlike in the image domain [14, 15], where masks can be readily obtained using additional models, the video domain incurs additional inference overhead and presents a more labor-intensive challenge due to the need for frame-by-frame mask generation. Recognizing the high redundancy between video frames, we propose a Point-Matching Mask Propagation strategy to efficiently extend the initial frame’s mask to subsequent frames. Leveraging feature maps from the inversion process [16], this strategy captures correlations in the latent space, eliminating the need for additional models.
@@
By applying the inversion on both the original video and style image, we obtain the initial noise for each. Integrated with our point-matching mask propagation strategy, we can also obtain per-frame masks for the object region specified by users. Then, we establish new edited latents and construct a three-branch denoising architecture with the content, edited, and style branches, for training-free AdaIN-Guided Localized Video Stylization. It first utilizes localized latent blending to fuse information from the content branch and edited branch, achieving localized style control during the style transfer process. In addition, to enhance information interaction between edited and style branches during denoising, we propose a training-free AdaIN-guided latent-shift and attention-shift mechanism. Unlike the image stylization method [17], It effectively mitigates detail loss while adeptly balancing content fidelity and style richness during the stylization.
@@ 2 Related Work
Figure 2: Overall framework of our UniVST. It consists of three core components. (1) Point-Matching Mask Propagation, (2) AdaIN-Guided Localized Video Stylization (Localized Latent Blending, AdaIN-Guided Latent-Shift and Attention-Shift) and (3) Sliding-Window Consistent Smoothing.
@@ 3 Preliminary
$Z_{t-1}=\sqrt{\alpha_{t-1}}\underbrace{{Z}_{t\rightarrow 0}}_{\text{predicted }Z_{0}}+\underbrace{\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}\left(Z_{t},t,C\right)}_{\text{direction pointing to }Z_{t-1}},$ (1)
@@
where $\mathcal{F}^{i}$ denotes the extracted feature of the $i$-th frame, while $W^{query}$, $W^{key}$, and $W^{value}$ represent the corresponding learnable mapping matrices, with [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] indicating concatenation operation.
@@ 4 Localized Video Style Transfer
7:Initialize: PreviousFeatures [$\leftarrow$[\,]← [ ] Initialize empty list for previous features.
8:Initialize: PreviousMasks [$\leftarrow$[\,]← [ ] Initialize empty list for previous masks.
9:Initialize: FirstFeature [$\leftarrow$[\,]← [ ] Initialize empty list for first feature.
@@
13:    $\text{ForeIndices}\leftarrow\operatorname{RandomSample}(\text{ForeIndices},r\cdot\frac{\lvert\text{ForeIndices}\rvert}{\lvert\text{ForeIndices}\rvert+\lvert\text{BackIndices}\rvert})$
14:    $\text{BackIndices}\leftarrow\operatorname{RandomSample}(\text{BackIndices},r\cdot\frac{\lvert\text{BackIndices}\rvert}{\lvert\text{ForeIndices}\rvert+\lvert\text{BackIndices}\rvert})$
@@
[…]
@@
4.1 Point-Matching Mask Propagation
“A towering building begins with a solid foundation.”
Compared to image style transfer, the video domain presents a unique challenge: obtaining mask information for each frame. One potential solution is to employ segmentation models, such as SAM [35], which can perform frame-by-frame segmentation based on the specified object. However, as shown in Fig. 3, this approach has notable drawbacks. Specifically, it introduces an additional memory overhead of 7.9 GB and is computationally intensive, which is unfavorable for deploying our UniVST framework.
@@ 4.1 Point-Matching Mask Propagation
In which, we can extend the initial frame’s mask across subsequent frames by the mask propagation strategy introduced below.
Mask Propagation. Given an initial object mask $\mathcal{M}^{1}$ specified by the users, we use $k$-NN [37] to propagate this mask to subsequent frames. For each point in the $i$-th frame, we identify the $k$ most similar points in the $(i-1)$-th frame based on the point-matching principle described above. If the majority of these $k$ points are in the foreground, the corresponding point in the $i$-th frame is classified as the foreground; otherwise, it is classified as the background. In this way, we can obtain the masks $\{\mathcal{M}^{i}\}_{i=1}^{N}$ for all frames.
@@ 4.2 AdaIN-Guided Localized Video Stylization
To achieve localized video stylization, we first separately obtain the initial latent of the original video $\{I^{i}\}_{i=1}^{N}$ and the style image $I^{s}$ using the inversion technique [38], denoted as $Z_{t}$ and $Z_{t}^{s}$ respectively, for $t=1\rightarrow T$. Integrated with our point-matching mask propagation strategy, we can also obtain the mask frame-by-frame efficiently. We then initialize the edited latent $\mathcal{Z}_{T}=Z_{T}$ and introduce a three-branch architecture with content, edited, and style branches, which perform the subsequent denoising process in parallel for $t=T\to 1$. These branches will interact through localized latent blending, AdaIN-guided latent-shift, and AdaIN-guided attention-shift, enabling precise control in subsequent video stylization.
@@
Figure 5: Comparison of training-free stylization methods: (a) original video frame, (b) key-value replacement [17], (c) key-value AdaIN and (d) AdaIN-guided attention-shift. Our approach mitigates local detail loss.
AdaIN-Guided Attention-Shift. In addition to transferring color tone, we also need to incorporate the overall style of the reference image into the edited video. The previous work StyleID [17] achieves style transfer by modifying the standard self-attention process. Specifically, the query $\mathcal{Q}_{t}$ in the editing branch attends to the key $K_{t}^{s}$ and value $V_{t}^{s}$ of the style branch instead of its own, referred to as K-V replacement. Furthermore, to eliminate content damage caused by direct replacement, [17] blends the original video’s query $Q_{t}$ with the edited query $\mathcal{Q}_{t}$. It is formulated as:
$\begin{split}\tilde{\mathcal{Q}}_{t}&=\gamma\cdot\mathcal{Q}_{t}+(1-\gamma)\cdot Q_{t},\\ &\tilde{\mathcal{K}}_{t}=K_{t}^{s},\;\tilde{\mathcal{V}}_{t}=V_{t}^{s},\\ &\operatorname{Attention}(\tilde{\mathcal{Q}}_{t},\tilde{\mathcal{K}}_{t},\tilde{\mathcal{V}}_{t}),\end{split}$ (9)
where $\gamma$ controls the balance between content and style. However, as observed in Fig. 5(b), K-V replacement still results in slight local artifacts in the stylized videos. To further reduce the local artifacts caused by direct replacement, we apply AdaIN to $\mathcal{K}_{t}$ and $\mathcal{V}_{t}$ in the editing branch, shifting their mean and standard deviation to align with $K^{s}$ and $V^{s}$ in the style branch, which can be described as:
@@
in which, $\beta_{\tau_{2}}$ and $\beta_{\tau_{3}}$ are the predefined weighting factors.
@@ 4.3 Sliding-Window Consistent Smoothing
Flickering and artifacts in the final edited video occur when AdaIN-guided localized video stylization is applied frame by frame. We empirically observe that this issue is caused by the training-free stylization mechanism, which disrupts the temporal consistency of the edited latent. Inspired by previous works [1, 2], we further incorporate optical flow with specialized sliding-window consistent smoothing to improve inter-frame coherence during denoising.
@@
Sliding-Window Smoothing. Rather than relying on simple post‐processing, we integrate the Warp operation into the denoising process to achieve local smoothing. At time step $t$, we calculate the latent representation $\mathcal{Z}_{t\rightarrow 0}$ from the predicted noise $\varepsilon_{t}$ according to Eq. (2). $\mathcal{Z}_{t\rightarrow 0}$ is then decoded by the decoder $\mathcal{D}$ into a pixel representation $\{P_{t}^{i}\}_{i=1}^{N}$. For each decoded frame $P_{t}^{i}$, we apply a window of size $2m$ to warp neighboring frames onto it, followed by averaging to smooth the original pixel of $P_{t}^{i}$, as detailed below:
@@
The above sliding-window consistent smoothing scheme is performed only within the interval $t\in[\tau_{4},\tau_{5}]$. Moreover, due to the challenges of optical flow estimation for fast-moving objects, we avoid global smoothing for long-range spatial alignment, opting instead for local sliding-window for short-range spatial alignment.
TABLE I: Quantitative comparison on DAVTG-WikiArt, where * denotes non-diffusion methods. The bold and underline indicate the best and second-best.
Methods Overall transfer performance Local style preservation Temporal consistency Inference efficiency
[…]
Diffutoon [12] 43.150 26.268 0.582 0.961 0.954 0.980 44.263
StyleCrafter [11] 39.119 21.008 0.777 0.928 0.895 0.978 111.225
AnyV2V [13] 33.774 19.104 0.680 0.938 0.914 0.978 78.474
BIVDiff [10] 46.071 25.468 0.740 0.921 0.894 0.955 158.367
EFDM [20] 37.407 22.350 0.602 0.936 0.935 0.963 5.969
CAST [19] 45.198 26.174 0.663 0.948 0.936 0.975 7.348
UniVST (ours) 37.152 26.134 0.369 0.986 0.991 0.980 162.745
TABLE II: Quantitative comparison on DAVTG-Laion, where * denotes non-diffusion methods. The bold and underline indicate the best and second-best.
Methods Overall transfer performance Local style preservation Temporal consistency Inference efficiency
[…]
Diffutoon [12] 38.610 23.700 0.563 0.959 0.953 0.979 44.263
StyleCrafter [11] 43.161 23.246 0.780 0.926 0.897 0.978 111.225
AnyV2V [13] 35.648 20.019 0.696 0.936 0.918 0.975 78.474
BIVDiff [10] 36.676 20.156 0.733 0.921 0.897 0.956 158.367
EFDM [20] 35.673 21.549 0.582 0.933 0.931 0.964 5.969
CAST [19] 34.828 19.974 0.661 0.947 0.936 0.968 7.348
UniVST (ours) 30.636 21.146 0.383 0.986 0.990 0.981 162.745
@@ 5.1 Experimental Settings
Our proposed UniVST is based on the Stable Diffusion V-1.5 [32]. The hyper-parameters are configured as follows: $\tau_{0}=0.1T$, $\tau_{1}=0.2T$, $\tau_{2}=0.4T$, $\tau_{3}=1.0T$, $\tau_{4}=0.5T$, $\tau_{5}=0.6T$, $t_{0}=0.4T$, with the total time period $T$ set to 50. Additionally, we set $\gamma=0.35$, $\beta_{\tau_{2}}=0.1$, $\beta_{\tau_{3}}=0.9$, $r=0.3$, $k=15$, $m=2$, and $n=9$. Original videos are resized to $512\times 512$ pixels and processed in batches of 16 frames. We use the RAFT model [41] for optical flow estimation during the denoising process. UniVST can be executed on a single NVIDIA GeForce RTX 3090 GPU without additional training or fine-tuning, making it both efficient and practical for deployment.

### v5

@@ 1 Introduction
Video stylization, a sub-area of video editing, is gaining popularity. It is about adding artistic styles to video content. Our review identifies two main video stylization methods. The first method uses the image or text to direct the style. StyleCrafter [11] personalizes video style with an extra training adapter; VideoBooth [8] uses the image for custom video creation prompts; while Diffutoon [12] focuses on cartoon coloring from a text prompt. The second method includes general video editing frameworks like AnyV2V [13] and BIVDiff [10] that can perform various tasks, including style transfer. These usually apply a consistent style across the whole video, which is good for thematic expression but sometimes lacks precision for specific uses. For example, as shown in Fig. 1, in film, scenes might be shot normally and later stylized to match different environments. This enables varied scene changes from one video take, which is also useful in advertising for efficiently applying local style changes to existing videos. Therefore, selectively styling parts of a video while keeping others unchanged is an important academic and commercial task.
@@
We illustrate Fig. 1 and pinpoint three major limitations of current methods: 1. Lack of Fine-grained Control. While some existing approaches [6, 5] leverage localized text descriptions to preserve the overall style transfer and maintain the primary object’s style, the control is usually too coarse. This lack of precision can lead to the model failing to understand the nuances, potentially causing unintended style transfer effects on objects not explicitly mentioned in the text. 2. Content Fidelity vs. Style Richness. In the realm of video stylization, there is a delicate balance to be struck between content fidelity and style richness. An overemphasis on style richness can result in blurring and altering the original layout of the video, whereas an excessive focus on content fidelity might yield a stylized outcome that is indistinguishable from the original. 3. Temporal Inconsistency. Unlike image stylization, video stylization requires careful consideration of temporal coherence between frames. As in BIVDiff [10] and AnyV2V [13], simply extending image stylization techniques to video often results in frame inconsistencies, causing various flickering and artifacts.
@@
In this paper, we propose UniVST, which integrates three components into a single training-free framework for the task of localized [^1] Here, “localized” indicates a particular part of the video. video style transfer, providing a unified solution to the above three limitations. The first step in achieving this is to obtain masks of the region specified by users, which allows for precise control during the subsequent stylization process. However, unlike in the image domain [14, 15], where masks can be readily obtained using additional models, the video domain incurs additional inference overhead and presents a more labor-intensive challenge due to the need for frame-by-frame mask generation. Recognizing the high redundancy between video frames, we propose a Point-Matching Mask Propagation strategy to efficiently extend the initial frame’s mask to subsequent frames. Leveraging feature maps from the inversion process [16], this strategy captures correlations in the latent space, eliminating the need for additional models.
@@
By applying the DDIM inversion on both the original video and style image, we obtain the initial noise for each. Integrated with our point-matching mask propagation strategy, we can also obtain frame-by-frame masks for the object region specified by users. Then, we establish new edited latents and construct a three-branch denoising architecture with content, edited, and style branches, for training-free AdaIN-Guided Localized Video Stylization. It first utilizes localized latent blending to fuse information from the content and edited branches, achieving localized style control during the style transfer process. In addition, to enhance information interaction between the edited and style branches during denoising, we propose a training-free AdaIN-guided latent-shift and attention-shift mechanism. Unlike traditional image stylization methods [17], It effectively mitigates detail loss during the stylization process, while also adeptly balancing content fidelity and style richness.
@@ 2 Related Work
Figure 2: Overall framework of our UniVST. It consists of three core components. (1) Point-Matching Mask Propagation, (2) AdaIN-Guided Localized Video Stylization (Localized Latent Blending, AdaIN-Guided Latent-Shift, and Attention-Shift), and (3) Sliding-Window Consistent Smoothing.
@@ 3 Preliminary
[…]
@@
where $\mathcal{F}^{i}$ denotes the extracted feature of the $i$-th frame, while $W^{query}$, $W^{key}$, and $W^{value}$ represent the corresponding learnable mapping matrices, with $\cdot$$\cdot$] indicating concatenation operation.
@@ 4 Localized Video Style Transfer
4.1 Point-Matching Mask Propagation
“A towering building begins with a solid foundation.”
Compared to image style transfer, the video domain presents a unique challenge: obtaining mask information for each frame. One potential solution is to employ segmentation models, such as SAM [35], which can perform frame-by-frame segmentation based on the specified object. However, as shown in Fig. 3, this approach has notable drawbacks. Specifically, employing SAM-H introduces an additional memory overhead of $7.9$ GB and increases the inference time, making it unfavorable for our UniVST.
@@ 4.1 Point-Matching Mask Propagation
7:Initialize: PreviousFeatures $\leftarrow[\,]$ Initialize empty list for previous features.
8:Initialize: PreviousMasks $\leftarrow[\,]$ Initialize empty list for previous masks.
9:Initialize: FirstFeature $\leftarrow[\,]$ Initialize empty list for first feature.
@@
[…]
@@
25:   $f\leftarrow\operatorname{Concat}(\text{PreviousFeatures},\text{FirstFeature})$
@@
In which we can extend the initial frame’s mask across subsequent frames by the mask propagation strategy introduced below.
Mask Propagation. Given an initial object mask $\mathcal{M}^{1}$ specified by the users, we use $k$-NN [37] to propagate this mask to subsequent frames. For each point in the $i$-th frame, we identify the $k$ most similar points in the (i1)(i-1)-th frame based on the point-matching principle described above. If the majority of these $k$ points are in the foreground, the corresponding point in the $i$-th frame is classified as the foreground; otherwise, it is classified as the background. In this way, we can obtain the masks $\{\mathcal{M}^{i}\}_{i=1}^{N}$ for all frames.
@@ 4.2 AdaIN-Guided Localized Video Stylization
To achieve localized video stylization, we first obtain the initial latent of the original video $\{I^{i}\}_{i=1}^{N}$ and the style image $I^{s}$ using the inversion technique [38], denoted as $Z_{t}$ and $Z_{t}^{s}$ respectively, for $t=1\rightarrow T$. Integrated with our point-matching mask propagation strategy, we can also obtain the mask frame-by-frame efficiently. We then initialize the edited latent $\mathcal{Z}_{T}=Z_{T}$ and introduce a three-branch architecture with content, edited, and style branches, which perform the subsequent denoising process in parallel for $t=T\to 1$. These branches will interact through localized latent blending, AdaIN-guided latent-shift, and AdaIN-guided attention-shift, enabling precise control in subsequent video stylization.
@@
Figure 5: Comparison of training-free stylization methods: (a) original video frame, (b) key-value replacement [17], (c) key-value AdaIN, and (d) AdaIN-guided attention-shift. Our approach mitigates local detail loss.
AdaIN-Guided Attention-Shift. In addition to transferring color tone, we also need to incorporate the overall style of the reference image into the edited video. The previous work StyleID [17] achieves style transfer by modifying the standard self-attention process, which is referred to as K-V replacement. Specifically, the query $\mathcal{Q}_{t}$ in the editing branch attends to the key $K_{t}^{s}$ and value $V_{t}^{s}$ of the style branch instead of its own. Furthermore, to eliminate content damage caused by direct replacement, [17] blends the original video’s query $Q_{t}$ with the edited query $\mathcal{Q}_{t}$. Finally, the resulting $\tilde{\mathcal{Q}}_{t}$, $\tilde{\mathcal{K}}_{t}$, and $\tilde{\mathcal{V}}_{t}$ undergo attention computation to produce the output $\mathcal{O}$. This entire process can be described as follows:
$\begin{split}\tilde{\mathcal{Q}}_{t}&=\gamma\cdot\mathcal{Q}_{t}+(1-\gamma)\cdot Q_{t},\\ &\tilde{\mathcal{K}}_{t}=K_{t}^{s},\;\tilde{\mathcal{V}}_{t}=V_{t}^{s},\\ &\mathcal{O}=\operatorname{Attention}(\tilde{\mathcal{Q}}_{t},\tilde{\mathcal{K}}_{t},\tilde{\mathcal{V}}_{t}),\end{split}$ (9)
where $\gamma$ controls the balance between content and style. However, as observed in Fig. 5(b), K-V replacement still results in slight local artifacts in the stylized videos. To further reduce the local artifacts caused by direct replacement, we apply AdaIN to $\mathcal{K}_{t}$ and $\mathcal{V}_{t}$ in the editing branch, shifting their mean and standard deviation to align with $K^{s}$ and $V^{s}$ in the style branch, resulting $\tilde{\mathcal{K}}_{t}$ and $\tilde{\mathcal{V}}_{t}$ as follows:
@@
in which $\beta_{\tau_{2}}$ and $\beta_{\tau_{3}}$ are the predefined weighting factors.
@@ 4.3 Sliding-Window Consistent Smoothing
Flickering and artifacts in the final edited video occur when AdaIN-guided localized video stylization is applied frame-by-frame. We empirically observe that this issue is caused by the training-free stylization mechanism, which disrupts the temporal consistency of the edited latent. Prior works like MV-Diffusion [39] leverage optical flow [1, 2, 39] to model trajectory and motion patterns, thereby enhancing temporal consistency in video generation. While this approach is also feasible for the stylization process, it requires retraining modules, which conflicts with our training-free design philosophy. By contrast, UniVST incorporates optical flow with specialized sliding-window consistent smoothing to improve inter-frame coherence during denoising in a training-free manner.
@@
Sliding-Window Smoothing. Rather than relying on simple post‐processing, we integrate the Warp operation into the denoising process to achieve local smoothing. At time step $t$, we calculate the latent representation $\mathcal{Z}_{t\rightarrow 0}$ from the predicted noise $\varepsilon_{t}$ according to Eq. (2). $\mathcal{Z}_{t\rightarrow 0}$ is then decoded by the decoder $\mathcal{D}$ into a pixel representation $\{P_{t}^{i}\}_{i=1}^{N}$. For each decoded frame $P_{t}^{i}$, we apply a window of size $2m$ to warp neighboring frames onto it, followed by averaging to smooth the original pixel of $P_{t}^{i}$, as follows:
@@
The above sliding-window consistent smoothing scheme is performed only within the interval $t\in[\tau_{4},\tau_{5}]$. Moreover, due to the challenges of optical flow estimation for fast-moving objects, we avoid global smoothing for long-range spatial alignment, employing instead a local sliding-window for short-range spatial alignment.
@@ 5.1 Experimental Settings
The main experiments of our proposed UniVST are based on the SD-v1.5 [32]. The hyper-parameters are configured as follows: $\tau_{0}=0.1T$, $\tau_{1}=0.2T$, $\tau_{2}=0.4T$, $\tau_{3}=1.0T$, $\tau_{4}=0.5T$, $\tau_{5}=0.6T$, $t_{0}=0.4T$, with the total time period $T$ set to 50. Additionally, we set $\gamma=0.35$, $\beta_{\tau_{2}}=0.1$, $\beta_{\tau_{3}}=0.9$, $r=0.3$, $k=15$, $m=2$, and $n=9$. Original videos are resized to $512\times 512$ pixels and processed in batches of $16$ frames. We use the RAFT model [42] for optical flow estimation during the denoising process. All main experiments can be executed on a single RTX 3090 GPU in a training-free manner. To further validate the generalization and scalability of the proposed UniVST, we extend it to four additional representative backbones, including SD-v2.1 [32], SD-v3.0 [43], SD-v3.5 [43], and AnimateDiff-v2 [44], with both quantitative and qualitative comparisons. Given that SD-v3.0 and SD-v3.5 are rectified-flow models, we employ RF-Solver [45], which is specifically designed for rectified flows, to replace the original DDIM inversion.
@@
TABLE I: Quantitative comparison on different datasets, where * denotes non-diffusion methods. The bold and underline indicate the best and second-best.
Methods Overall transfer performance Local style preservation Consistency Efficiency Overall Quality
[…]
DAVTG-WikiArt
Diffutoon [12] 43.150 26.268 0.582 0.961 0.954 0.980 44.263 72.251
StyleCrafter [11] 39.119 21.008 0.777 0.928 0.895 0.978 111.225 70.759
AnyV2V [13] 33.774 19.104 0.680 0.938 0.914 0.978 78.474 72.791
BIVDiff [10] 46.071 25.468 0.740 0.921 0.894 0.955 158.367 70.351
EFDM [20] 37.407 22.350 0.602 0.936 0.935 0.963 5.969 71.649
CAST [19] 45.198 26.174 0.663 0.948 0.936 0.975 7.348 70.429
UniVST (SD-v1.5) 37.152 26.134 0.369 0.986 0.991 0.980 162.745 74.014
DAVTG-Laion
Diffutoon [12] 38.610 23.700 0.563 0.959 0.953 0.979 44.263 73.607
StyleCrafter [11] 43.161 23.246 0.780 0.926 0.897 0.978 111.225 71.379
AnyV2V [13] 35.648 20.019 0.696 0.936 0.918 0.975 78.474 71.927
BIVDiff [10] 36.676 20.156 0.733 0.921 0.897 0.956 158.367 70.273
EFDM [20] 35.673 21.549 0.582 0.933 0.931 0.964 5.969 70.628
CAST [19] 34.828 19.974 0.661 0.947 0.936 0.968 7.348 71.906
UniVST (SD-v1.5) 30.636 21.146 0.383 0.986 0.990 0.981 162.745 74.495